# Quantifying individual and team performance in cricket

### v1

@@ Abstract
Advancements in technology have recently allowed to collect and analyze large-scale fine-grained data about human performance, drastically changing the way we approach sports. Here we provide the first comprehensive analysis of individual and team performance in cricket, one of the most popular sports in the world. We investigate temporal patterns of individual success, finding that the best performance of a player can happen at any time in their career, surrounded by a burst of comparable top performances. We show that long-term performance can be predicted from early observations and that temporary exclusions of players from teams are often due to declining performances but are also associated with strong comebacks. We demonstrate that teams are more likely to endure hot streaks of winning performances compared to random expectations, but also sequences of consecutive losses. We investigate the crucial contributions of specialists such as openers, all-rounders and wicket-keepers and show that a balanced performance from multiple individuals is required to ensure team success. Finally, we quantify how transitioning to captaincy in the team improves the performance of batsmen, but not that of bowlers. Our work emphasizes how individual endeavours and team dynamics interconnect and influence collective outcomes in sports.
@@ I Introduction
The origin of sports can be traced back to early civilizations, notably with the inception of the Olympic Games in ancient Greece Roche (2000); Guttmann (2002); Bryson (2008); Horne and Whannel (2020). These early sporting events were not just about physical achievements but also played a role in cultural and societal bonding. As societies progressed, sports evolved, reflecting changes in social structures and becoming more organized and diverse. The recent advancements in digital technology have significantly altered the landscape of sports, introducing a new era of sports analytics Nevill et al. (2008). Enhanced data acquisition capabilities have resulted in large-scale data sets which lend themselves to quantitative analysis of athletes and teams. This evolution has not only heightened the excitement for fans but also provided researchers and sports strategists with valuable insights into the physical and cognitive skills essential in competitive sports Radicchi (2011); Pappalardo et al. (2019); Chowdhary et al. (2023); Zappalà et al. (2023).
In the realm of baseball, the influence of sports analytics has been particularly noteworthy, a phenomenon popularized by the ‘Moneyball’ revolution Lewis (2004). The strategic use of data analytics in baseball has led to a profound shift in team management and player evaluation, where data-driven decisions have often trumped traditional scouting methods Brown et al. (2017); Pappalardo and Cintia (2018). NBA and NFL have also embraced this analytical wave and now leverage data to optimize player performance, team strategies, and even in-game decisions Mason and Foster (2007). In the NBA, for instance, analytics have led to an increased emphasis on three-point shooting and defensive metrics, reshaping the style of play Dawson (2023); Neuhaus and Thomas (2024). On the other hand, more intellectual games like chess are also seeing an elevated level of sophistication and achievements owing to the introduction of advanced chess engines Roring and Charness (2007); Vaci and Bilalić (2017).
@@
Cricket, the world’s second most-followed sport after soccer according to multiple sources noa (2023, 2024), enjoys widespread popularity primarily in Commonwealth countries, including India, Australia, New Zealand, the United Kingdom, South Africa, the West Indies, Sri Lanka, and Pakistan. The availability of match data, driven by amateur and professional enthusiasts, has fostered various analyses, ranging from quantifying the hot-hand effect Ram et al. (2022) to predicting match outcomes using machine learning Subasingha et al. (2019), assessing cricket-related injuries J Orchard et al. (2002); Finch et al. (1999); Orchard et al. (2005), and modelling the game with state-of-the-art techniques Swartz et al. (2009); Kimber and Hansford (1993); Arora et al. (2014); K. (1905); Mehta et al. (1983). Despite this, there remains a substantial gap in the understanding of individual performances and their contributions to team success in cricket. In this work, we track players’ careers, unveiling universal patterns of performance and identifying correlates of team-level success.
We categorize players into batsmen and bowlers, adjusting their performance metrics for inflation, and examine the emergence of phenomena such as the random impact rule and hot streaks. These phenomena, extensively studied in diverse fields like science Sinatra et al. (2016); Liu et al. (2018, 2022), arts Fraiberger et al. (2018); Janosov et al. (2020); Williams et al. (2019), and sports Bar-Eli et al. (2006); Gilovich et al. (1985); Chowdhary et al. (2023), provide a framework for understanding patterns in cricket. Additionally, we investigate performance predictors, finding that early career achievements are indicative of overall career trajectories. We also explore factors contributing to player exclusion from teams. Building on these insights we assess the impact of leadership roles and pathways to captaincy and analyze the roles and contributions of specialists, such as openers, all-rounders, and wicket-keepers. Lastly, we apply established metrics from the literature to quantify patterns in team success. Collectively, our work presents the first comprehensive analysis of individual and team performance in cricket.
Figure 1: Time of top performance in a player’s career. (a) Distribution of the match number ($N^{*}$) of the top performance for a player’s career. The inset shows the distribution of the time of the top performance in a career $N^{*}$ normalized by the career length $N$. Real data (pink circles) is consistent with a randomized null model which removes temporal correlations (black dotted line) suggesting that the best performance can occur at any time within the career. (b) Distribution of differences between the locations of the best ($N^{*}$), second best ($N^{**}$), and third best ($N^{***}$) performances of a player normalized by their career lengths. Higher values near zero compared to the null-model indicate that the best performances are clustered, indicating the presence of hot streaks in cricket careers.
II Results
Data
@@ II Results
In this work, we conduct a comprehensive analysis of a dataset comprising approximately 4,500 One-Day International (ODI) cricket matches, extracted from howstat.com, an open-access repository for cricket statistics. This dataset encompasses records of all ODI matches played from their inception in 1971 until October 2023 (see Methods). Our data extraction focused on key metrics, including individual player performance (quantified in terms of runs scored and wickets taken), team scores, match winners, venues, and match dates. Although the primary aggregation of data is at the team level for each match, the dataset is also suitable for an in-depth exploration of individual player trajectories and careers. We have 607 batsmen and 594 bowlers in our dataset (see Methods). This allows for a multifaceted analysis of performance trends, the impact of various factors on player and team success, and the evolution of the sport over more than five decades.
@@
Team scores per match have generally increased over the decades (see SM1). This observed variation is potentially due to a confluence of factors, including modifications in game regulations (fielding rules for example), advancements or changes in the equipment used (notably the cricket bat), and the reduction in the dimensions of the playing field 65-75 metres compared to 80-85 metres in stadiums earlier as spectators are interested in high scoring matches. To facilitate a robust and equitable comparison of player and team performances across distinct time periods, we account for the inflation in run-scoring by implementing a normalization procedure on all batting statistics. Specifically, we adjusted these statistics such that the average team score is constant over the examined time period, following a well-known procedure originally introduced to correctly assess the impact of scientific papers published in different years Radicchi et al. (2008) (see SM1). This approach ensures that the comparative analysis of players’ performance from different eras is conducted in an unbiased manner which controls for background temporal trends.
II.1 Temporal patterns of top performances
@@
The progression of a player’s skill level enhances their likelihood of surpassing previous performance peaks. Conversely, as players age, their athletic prowess may diminish, potentially impeding their ability to exceed past achievements. This contrast of skill development against physical decline poses a critical question: at which point in their careers do players deliver their best performance? We address this question by making use of tools and methodologies from data science and extreme value theory already deployed in the realms of arts, culture, and more recently in sports Coles (2001); Beirlant et al. (2004); Sinatra et al. (2016); Janosov et al. (2020); Chowdhary et al. (2023). We define $N^{*}$ as the match number where a player achieves their career-best performance. For batsmen, this is marked as their highest run score. For bowlers, it corresponds to the highest number of wickets taken. In instances of identical runs or wickets, the performance involving respectively fewer balls played or fewer runs conceded is considered. The probability distribution function $P(N^{*})$ for all players is shown in fig. (1)(a). We observe that $P(N^{*})$ is exponentially distributed, suggesting a much higher likelihood of peak performance occurring earlier rather than later in a career. However, this analysis does not account for variations in career lengths. When normalizing the timing of peak performance relative to overall career length, we observe a uniform distribution, as shown in fig. (1)(a) (inset). This pattern, previously dubbed as the ‘random impact rule’ for scientific careers Sinatra et al. (2016), suggests that the timing of peak performance is generally unpredictable and can occur at any point in a career. Indeed, Sachin Tendulkar, one of the most prolific batsmen across generations has $N^{*}/N=0.95$, but for Virat Kohli $N^{*}/N=0.3$. Similarly among the bowlers, Muttiah Muralitharan, the highest wicket-taker in ODI has $N^{*}/N=0.43$, but Waqar Younis has $N^{*}/N=0.81$. To assess the robustness of our results, we compared them against a randomized null model which shuffles the order of performances for each player (see Methods), denoted by a dotted line in fig. (1). We find no qualitative difference between the empirical and randomized data distributions.
Across a range of domains such as science Liu et al. (2018, 2022), arts Williams et al. (2019); Janosov et al. (2020), and sports Ram et al. (2022); Chowdhary et al. (2023), individual careers display hot streaks, where exceptional performances tend to occur in bursts which are clustered in time. To extend this inquiry to the realm of cricket, we designate $N^{*}$, $N^{**}$, and $N^{***}$ as the match numbers corresponding to a player’s best, second-best, and third-best performances, respectively. A small temporal gap between $N^{*}$, $N^{**}$, and $N^{***}$ would indicate a concentration of top performances within a specific timeframe, indicating the existence of a hot streak. Conversely, a larger difference among these time stamps would suggest a lack of correlation in the timing of peak performances. For e.g. Rohit Sharma, who holds the record for the highest individual score in a match achieved this feat in his $121^{st}$ match ($N^{*}$=121). However, his second-highest score came in his $105^{th}$ match ($N^{**}=105$, $|N^{*}-N^{**}|/N=0.06$. In fig. (1)(b) we plot the distribution of the normalized number of matches between the best and second-best performances, as well as between the best and third-best performances. Additionally, we employ a null model (same as before) where we reshuffle the timestamps of performances and analyze the corresponding differences to benchmark our findings. Our analysis reveals that the probability distributions of $|N^{*}-N^{**}|$ and $|N^{*}-N^{***}|$ are consistently higher around smaller values ($\approx 0$) than those derived from the randomized null model. Moreover, the probability of performing well in consecutive matches grows exponentially as evidenced by the recurrence time distribution (see SM2).
In conclusion, while the best performance of a cricket player’s career may manifest at any point within their career, these further observations indicate that peak performances in cricket tend to occur in bursts over short periods of time.
II.2 Long-term performance from early career observations
Early identification of talent and skills can provide key advantages in many competitive settings, from firm growths Golder and Tellis (1997) to information spreading Bak-Coleman et al. (2022). Also in sports, nurturing talent in young players can lead to higher returns Baker et al. (2018); Till and Baker (2020); Zappalà et al. (2024). Here we investigate the relationship between a player’s initial performance and their overall career trajectory to capture whether it is possible to see hints of future performance based on early-career display.
In figures (2) (a) and (b) we conduct an analysis correlating early and overall career statistics. Our analysis considers players who have participated in a minimum of 50 matches. We focus on their mean performance during the initial 25 matches. For batsmen, we look at the average runs scored per match in this early phase, whereas for bowlers, we look at the average number of wickets taken per match within the same period. Concurrently, we examine the full-career performance of these players. The scatter plots in figure (2) reveal a correlation between early and full career average performance metrics. We calculate the linear regression coefficients $R^{2}=0.49$ for (a) batsmen and $R^{2}=0.66$ for (b) bowlers. This finding suggests that a player’s initial performance may serve as an indicator of their subsequent career performance. However, notable differences emerge when comparing batsmen and bowlers. For illustration, consider Ricky Ponting (batsman) and Shane Warne (bowler) from the Australian team. Ricky Ponting averaged 32.45 runs per match in his early career and later improved his career average to 37.5 runs per match. In contrast, Shane Warne picked up 2.12 wickets per match in his early career, but ultimately his performance dropped to 1.53 wickets per match over his full career. In general, approximately 55% of batsmen are observed to improve upon their early career averages, whereas this figure drops to about 45% for bowlers. Such disparities may be attributed to the distinct elements inherent in the roles of batsmen and bowlers. Batsmen, for instance, often leverage accumulated experience and refined skills to augment run-scoring, while bowlers may depend more on innovation and strategic creativity to increase their wicket tally. This distinction potentially accounts for the observed trend of a higher proportion of batsmen surpassing their initial career performance compared to bowlers. Alternatively, our observations can also be explained if some players are simply better than the rest. Thus, their performance in various stages of their careers will correlate with their full career performance. This effect is also known as the Q-model in the context of scientific careers Sinatra et al. (2016). Further analyses supporting this hypothesis are presented in SM3.

### v2

# Individual and team performance in cricket

@@ Abstract
Advancements in technology have recently allowed to collect and analyse large-scale fine-grained data about human performance, drastically changing the way we approach sports. Here we provide the first comprehensive analysis of individual and team performance in One-Day International cricket, one of the most popular sports in the world. We investigate temporal patterns of individual success by quantifying the location of the best performance of a player and find that they can happen at any time in their career, surrounded by a burst of comparable top performances. Our analysis shows that long-term performance can be predicted from early observations and that temporary exclusions of players from teams are often due to declining performances but are also associated with strong comebacks. By computing the duration of streaks of winning performances compared to random expectations, we demonstrate that teams win and lose matches consecutively. We define the contributions of specialists such as openers, all-rounders, and wicket-keepers and show that a balanced performance from multiple individuals is required to ensure team success. Finally, we measure how transitioning to captaincy in the team improves the performance of batsmen, but not that of bowlers. Our work emphasises how individual endeavours and team dynamics interconnect and influence collective outcomes in sports.
@@ I Introduction
The inception of sports, notably the Olympic Games in ancient Greece, played a pivotal role in cultural and societal bonding Roche (2000); Guttmann (2002). As societies evolved, sports mirrored changes in social structures, becoming more organised and diverse Bryson (2008); Horne and Whannel (2020). Recent digital technology advancements and enhanced data acquisition capabilities have ushered a new era of sports analytics, providing valuable insights into athlete and team performance Nevill et al. (2008); Radicchi (2011); Pappalardo et al. (2019); Chowdhary et al. (2023); Zappalà et al. (2023). In baseball, the ‘Moneyball’ revolution popularised the strategic use of data analytics, profoundly altering team management and player evaluation Lewis (2004); Brown et al. (2017); Pappalardo and Cintia (2018). Premier leagues in other sports such as basketball (NBA) and American football (NFL) have similarly embraced analytics to optimise player performance, team strategies, and in-game decisions, leading to stylistic shifts in play Mason and Foster (2007); Dawson (2023); Neuhaus and Thomas (2024). Intellectual games like chess have also advanced with the introduction of sophisticated chess engines Roring and Charness (2007); Vaci and Bilalić (2017); Chowdhary et al. (2023).
@@
One-Day International (ODI) Cricket, the world’s second most-followed sport after soccer noa (2023, 2024), enjoys widespread popularity primarily in Commonwealth countries, including India, Australia, New Zealand, the United Kingdom, South Africa, the West Indies, Sri Lanka, and Pakistan. The availability of match data, driven by amateur and professional enthusiasts, has fostered various analyses. One of the major lines of research has been predicting the match outcomes of ODI matches using a variety of techniques such as machine learning Subasingha et al. (2019), logistic regression using pre-match covariates Bandulasiri (2008); McEwan et al. (2023), and logistic regression using in-game dynamic variables Asif and McHale (2016). Other studies have tried to uncover specific patterns based on performance such as the hot-hand effect Ram et al. (2022) and ranking of the players Premkumar et al. (2020) or model the dynamics of the game Swartz et al. (2009). A few studies have tried to investigate in detail the batting K. (1905); Kimber and Hansford (1993) and bowling Mehta et al. (1983) aspects of the game. With the advent of shorter and faster formats of the game such as T20, some attention has been devoted to investigating the effect of premier leagues on social media activity Arora et al. (2014) as well as on other formats of the game Nicholls et al. (2023). Going beyond the specificities of ODI, researchers have tried to examine the role of injuries in cricket J Orchard et al. (2002); Finch et al. (1999); Orchard et al. (2005). Despite this, there remains a substantial gap in the understanding of individual performances and their contributions to team success in One-day international cricket. In this work, we track players’ careers, unveiling universal patterns of performance and identifying correlates of team-level success.
The progression of a player’s skill level enhances their likelihood of surpassing previous performance peaks Ram et al. (2022). Conversely, as players age, their athletic prowess may diminish, potentially impeding their ability to exceed past achievements J Orchard et al. (2002). This contrast of skill development against physical decline poses a critical question: at which point in their careers do players deliver their best performance? We address this question in Sec.(III.1) by making use of tools and methodologies from data science and extreme value theory Coles (2001); Beirlant et al. (2004) already deployed in diverse fields like science of science Sinatra et al. (2016); Liu et al. (2018, 2022), arts Fraiberger et al. (2018); Janosov et al. (2020); Williams et al. (2019), and sports Bar-Eli et al. (2006); Gilovich et al. (1985); Chowdhary et al. (2023).
Early identification of talent and skills can provide key advantages in many competitive settings, from firm growths Golder and Tellis (1997), information spreading Bak-Coleman et al. (2022) to sports, where nurturing talent in young players can lead to higher returns Baker et al. (2018); Till and Baker (2020); Zappalà et al. (2024). In Sec. (III.2) we extend this inquiry to investigate the relationship between a player’s initial performance and their overall career trajectory to capture whether it is possible to see hints of future performance based on early-career display.
Fluctuations in team composition frequently arise due to injuries and variations in player performance Pasarakonda et al. (2023). While injuries often occur unpredictably, a decline in performance typically manifests more gradually and may not be immediately apparent. A fitting inquiry in this context is whether it is possible to discern patterns in a player’s performance preceding their expulsion from the team. We study this aspect in Sec. (III.3).
The composition of an effective team encompasses skilled players under proficient leadership. Case studies across various domains, including sports, science, and business, have illustrated scenarios, where moderately weak teams achieve success under adverse conditions, guided by strategic leadership Guimerà et al. (2005); Zhang et al. (2022); Hancock et al. (2023); Betti et al. (2024). Here, we ask the reverse question, is a player’s performance affected by the burden of leadership in Sec. (III.4). However, from a collective perspective, the strategic composition of a team is crucial for its effective functioning. The concept of utilising specialists – individuals who perform highly specific roles – extends beyond sports into various domains of human activity, including business organisations, scientific research, software development, and even hunter-gatherer societies Burke et al. (2019); Salcinovic et al. (2022); Wallrich et al. (2024). We analyse the role of three specialists – openers, all-rounders, and wicket-keepers in Sec. (III.5).
So far we have predominantly focused on individual contributions to team success in ODI cricket. We now broaden our perspective to analyse team success, setting aside specific individualistic factors that contribute to victory. Given that each cricket match culminates in a definitive outcome, our interest lies in discerning patterns of wins and losses for teams. We apply established metrics from the literature to quantify patterns in team success in Sec. (III.6). Collectively, our work presents the first comprehensive analysis of individual and team performance in ODI cricket.
II Methods and Materials
II.1 Data collection
We extracted data on 4418 One Day International (ODI) matches in men’s cricket, involving 2863 unique players by web scrapping howstat.com (Ref. how (2024)), an open-access repository for cricket statistics, using BeautifulSoup and urllib, two python libraries. This dataset encompasses records of all ODI matches played from their inception in 1971 until March 2024. For each match, we extracted information including date, teams, runs scored, wickets taken, and overs played by each team, along with player names and their contributions in terms of batting and bowling. This included details like batting position, number of balls faced, number of fours and sixes hit, number of overs bowled, number of runs conceded, number of maiden overs, and number of wickets taken. Furthermore, data on the captain and wicket-keeper for each match were also collected.
@@ II.1 Data collection
Although the primary aggregation of data is at the team level for each match, the dataset is also suitable for an in-depth exploration of individual player trajectories and careers. This allows for a multifaceted analysis of performance trends, the impact of various factors on player and team success, and the evolution of the sport over more than five decades.
II.2 Classification of players
We give a brief description of our methodology to classify players. For curious readers, we direct to Ref. cri (2024) for the exact role of each type of player. Note that a player can be simultaneously classified into multiple categories. For e.g., a batsman can be an all-rounder, a captain, an opener, and a wicket-keeper.
Batsman: We consider players as batsmen who have played at least 25 matches and batted at position 7 or above in at least 50% of their matches. We have 580 batsmen in our dataset after applying this criteria.
Bowlers: We classify players as bowlers who have played at least 25 matches and bowled in at least 50% of the matches they played. We have 551 bowlers in our dataset after applying this criterion.
All-rounders: We consider all-rounders as players who have played at least 25 matches and are classified as both batsmen and bowlers using the above definitions. We have 117 all-rounders in our dataset.
Captains: In our analysis, we consider players as captains if they have captained the team in at least 15 matches. The information about captaincy is available as metadata on the website.
Openers and Non-openers: We characterise openers as players who have batted at first or second position at least 10 times in their career, while non-openers as players who have batted at positions three to six at least 10 times in their career.
Wicket-keepers: In our analysis, we consider players as wicket-keepers based on the metadata available for each match on the website
Fielders: All players on the bowling side who are not bowling or wicket-keeping are classified as fielder in our analysis.
II.3 Data normalisation
@@
Team scores per match have generally increased over the decades (see SM1). This observed variation is potentially due to a confluence of factors, including modifications in-game regulations (fielding rules for example)Smyth (2011), advancements or changes in the equipment used (notably the cricket bat) Sankar (2019), and the reduction in the dimensions of the playing field 65-75 metres compared to 80-85 metres in stadiums earlier as spectators are interested in high scoring matches Silgardo (2023). To facilitate a robust and equitable comparison of player and team performances across distinct time periods, we account for the inflation in run-scoring by implementing a normalisation procedure on all batting statistics. Specifically, we multiplied runs scored by the batsmen and the runs given by a bowler in a given year by a normalisation factor $nf=\frac{\langle Team\;runs\rangle_{all}}{\langle Team\;runs\rangle_{year}}$ such that the average team score is constant over the examined time period. Here, $\langle Team\;runs\rangle_{all}$ is the average team score across all years, while $\langle Team\;runs\rangle_{year}$ is the average team score in the given year. This procedure was originally introduced to correctly assess the impact of scientific papers published in different years Radicchi et al. (2008). Our approach ensures that the comparative analysis of players’ performance from different eras is conducted in an unbiased manner that controls for background temporal trends.
II.4 Statistical data analysis
II.4.1 Statistical tests
Linear regression is used to estimate the linear relationship between two variables and returns a correlation coefficient. We use it in Sec. (III.2).
Kolmogorov-Smirnov test (K-S test for short) is a non-parametric test used to determine if two unpaired sampled distributions come from the same underlying distribution Hodges (1958). The null hypothesis is that the two distributions are the same and the $p$-value gives the probability that the samples in question are taken from the same distribution. We use the K-S test in Secs. (III.1) and (III.5).
Wilcoxon signed rank test is a non-parametric test used to determine if the location of means of two paired distributions is the same Wilcoxon (1945). Assuming the null hypothesis that the means are the same, it returns the probability that the null hypothesis is true. We use the Wilcoxon signed rank test in Sec. (III.1).
Mann-Whitney $U$ test is a non-parametric test of the null hypothesis that the unpaired distributions underlying the two samples are the same Mann and Whitney (1947). This test checks the null hypothesis if one of the distributions is stochastically larger or smaller than the other. We use the Mann-Whitney $U$ test in secs. (III.4), (III.5), and (III.6).
Welch’s t-test is used to determine if the two populations have similar means when the variances and sample sizes may or may not be equal Welch (1947). We use this test in Sec. (III.5).
All the above tests are conducted using the SciPy package of Python Virtanen et al. (2020).
II.4.2 Fractional contribution
To quantitatively and equitably assess the impact of each player, we introduce the fractional contribution $f_{c}$ as,
, so that a player’s maximum possible contribution ($f_{c}$) is normalised within the range of 0 to 1. It is important to note that, given this framework, specialised batsmen and bowlers have an upper limit of $f_{c}=0.5$ in a best-case scenario.We use this definition in Secs. (III.5) and (III.6).
II.4.3 Effective team size
We employ the concept of effective team size, a well-established metric in the team analysis literature Klug and Bagrow (2016); Delice et al. (2019) to characterise the heterogeneous contributions of individuals to the team. Effective team size quantifies the essential number of players in a team, effectively measuring the redundancy in team composition. Mathematically, we denote the contribution of the player to the team score as $f_{c}$, such that $\sum_{c}f_{c}=1$, as per Eq. (1). The effective team size $S_{eff}$ is defined in such a way that it equals the actual team size if contributions are evenly distributed among all players and reduces to 1 if a single player is solely responsible for the entire performance. A common approach to calculate $S_{eff}$ is to use the formula $S_{eff}=2^{H}$, where $H=-\sum_{c}f_{c}\log_{2}f_{c}$. Statistically, $S_{eff}$ represents the entropy of the distribution of contributions $f_{i}$ from team players. We use this concept in Sec. (III.6).
II.5 Null models
Figure 1: Time of top performance in a player’s career. (a) Distribution of the match number ($N^{*}$) of the top performance for a player’s career. The inset shows the distribution of the time of the top performance in a career $N^{*}$ normalised by the career length $N$. Real data (pink circles) is consistent with a randomised null model which removes temporal correlations (black dotted line) suggesting that the best performance can occur at any time within the career. (b) Ratio of number of players for the original data ($n^{data}_{\Delta N}$) and randomised data ($n^{rnd}_{\Delta N}$) having normalised differences ($\Delta N/N$) between best performances ($|N^{*}-N^{**}|$ in blue and $|N^{*}-N^{***}|$ in brown) in each particular bin. Values higher than 1 for small $\Delta N/N$ indicate that the best performances are clustered, highlighting the presence of hot streaks in cricket careers.
In our work, we analyse characteristic patterns in the performance of both players and teams. In order to control for confounding variables, one of the usual approaches is to employ null models. Null model is a concept commonly used to capture the existence (or lack thereof) of signal from noisy data. It often involves randomizing certain aspects of the data while keeping the rest of the factors constant. Comparing the differences between the randomised and true data allows us to extract statistically significant and meaningful patterns. Here, we use two null models for detecting individual and team performance signatures.
II.5.1 Best individual performance
In Sec. (III.1), we characterise the temporal patterns of top performances. In order to establish a baseline for the measures, we shuffle the timestamps of individual performances for each player 100 times keeping the actual values of performance the same. In this way, we are breaking the temporal correlation between different performances. For each shuffling, we find the time stamp for the best ($N^{*}$), second best ($N^{**}$), and third best ($N^{***}$) performances of a player. We compare this randomly shuffled data to the original dataset to establish the similarities and differences.
II.5.2 Hot and cold streaks for teams
In Sec. (III.6) we show the existence of hot and cold streaks for teams. We consider all the teams that have played at least 200 matches. By keeping the number of wins constant, we shuffle the timestamps of the match results $10^{4}$ times for each team. For each reshuffling, we compute the number of consecutive wins (hot) and losses (cold) denoted by $L_{streak}$. By taking the average across $10^{4}$ reshuffles, we compute $N_{rand}$ to highlight the significance of $N_{data}$ – the number of streaks of length $l$ or higher in the actual dataset.
III Results
III.1 Temporal patterns of top performances
@@
Methods: We designate $N^{*}$, $N^{**}$, and $N^{***}$ as the match numbers corresponding to a player’s best, second-best, and third-best performances, respectively. For batsmen, this is marked as their highest run score. For bowlers, it corresponds to the highest number of wickets taken. In instances of identical runs or wickets, the performance involving respectively fewer balls played or fewer runs conceded is considered. Additionally, to account for variations in the career lengths of the individual players, we normalise the timing of peak performances ($N^{*}$, $N^{**}$, and $N^{***}$) with overall career lengths ($N$).
We are interested in two questions, (1) Does the best performance ($N^{*}$) occur at a specific time within a player’s career? and (2) Are the best performances closely related in time? To answer the first question, we calculate the probability distribution function of $N^{*}$ as well as $N^{*}/N$ and compare it with the randomised dataset as shown in Sec. (II.5). This analysis was originally introduced in Ref. Liu et al. (2018) to test the presence of hot-streaks in time series data. We use the K-S test to check if the distributions obtained from the original and randomised datasets differ. On the other hand, to investigate the second question, we divide the differences between the top performances $\Delta N$, i.e. $|N^{*}-N^{**}|$ and $|N^{*}-N^{***}|$, normalised by the career length ($N$) into 5 bins. Within each bin, we compute the ratios of the number of players in the data ($n^{data}_{\Delta N}$) and the number of players in the randomised data ($n^{rnd}_{\Delta N}$). We run the Wilcoxon signed rank test to determine if the values of the ratios differing from 1 are significant.
Results: The probability distribution function $P(N^{*})$ for all players is shown in fig. (1)(a). We observe that $P(N^{*})$ is monotonically decreasing function, suggesting a much higher likelihood of peak performance occurring earlier rather than later in a career. However, this analysis does not account for variations in career lengths. When normalising the timing of peak performance relative to overall career length, we observe a uniform distribution, as shown in fig. (1)(a) (inset). This pattern, previously dubbed as the ‘random impact rule’ for scientific careers Sinatra et al. (2016), suggests that the timing of peak performance is generally unpredictable and can occur at any point in a career. The K-S test comparing the data to the null-model gives $p$-value $>0.05$ in both cases, thus giving an indication that the null hypothesis of the two distributions being statistically the same can not be rejected.
In fig. (1)(b) we plot the ratios of the number of players in the bins of $\Delta N/N$ for the original and the randomised dataset. We notice that, for $\Delta N/N\in[0.0,0.2)$ (indicating top performances closely related in time) we observe a ratio $>1$, signifying that the number of players having small gaps in their best performances is higher than expected by chance on a randomised data. In other words, they exhibit hot-streaks – brief periods of time accentuated by best performances of comparable magnitude. In contrast, the ratio of the number of players for $\Delta N/N\in[0.6,1.0]$ (indicating that best performances are far away in time) is always less than one. This implies that fewer players have long gaps between best performances compared to random expectations. The Wilcoxon signed rank test indicates a significant deviation of ratios from 1 ($p$-value $<0.001$) for $\Delta N/N\in[0.0,0.2)$ and $\Delta N/N\in[0.6,1.0]$ for both $|N^{*}-N^{**}|$ and $|N^{*}-N^{***}|$. Our results are robust to the number of bins as seen in SM2.
III.2 Long-term performance from early career observations
We are interested in determining the relationship between a player’s early career trajectory and their overall career performance. Methods: In figures (2) (a) and (b) we correlate early and overall career statistics. Our analysis considers players who have participated in a minimum of 50 matches. We focus on their mean performance during the initial 25 matches. For batsmen, we look at the average runs scored per match in this early phase, whereas for bowlers, we look at the average number of wickets taken per match within the same period. Concurrently, we examine the full-career performance of these players.
Results: The scatter plots in figure (2) reveal a correlation between early and full career average performance metrics. We calculate the linear regression coefficients $R^{2}=0.45$ for (a) batsmen and $R^{2}=0.66$ for (b) bowlers. This finding suggests that a player’s initial performance may serve as an indicator of their subsequent career performance. However, notable differences emerge when comparing batsmen and bowlers. In general, approximately 55% of batsmen are observed to improve upon their early career averages, whereas this figure drops to about 45% for bowlers.